# Data-driven controlled subgroup selection in clinical trials

### v1

@@ 3.3 Subgroup selection for heterogeneous treatment effect
Now, if in addition to the observed response $Y_{i}(T_{i})$, we would observe the counterfactual response $Y_{i}(1-T_{i})$, we could take $Y_{i}(1)-Y_{i}(0)$ as the input for the response in the methods described in the previous subsections, as the covariate-conditional expectation of that difference is exactly the CATE, so that the output of the above methods would give an estimate of $S_{\tau}$. Unfortunately, this is not possible in practice.
The idea of the meta-learning framework is then to use flexible machine learning methods in order to estimate nuisance functions like the propensity $\pi:x\mapsto\mathbb{P}(T_{i}=1\mid X_{i}=x)$ and the treatment-conditional regression functions $\eta_{t}(x):=\mathbb{E}\bigl\{Y_{i}(T_{i})\mid T_{i}=t,X_{i}=x\bigr\}$ for $t\in\{0,1\}$. Following the taxonomy of curth2021nonparametric, one-step plug-in learners would then estimate the CATE directly from the estimates of $\eta_{t}$, $t\in\{0,1\}$, whereas two-step learners use these estimates in order to generate pseudo-observations serving as proxies for the unobservable differences $Y_{i}(1)-Y_{i}(0)$. In particular, one may generate these pseudo-observations using the double robust (DR) learning method described by kennedy2023optimal. For that purpose, one calculates estimators of $\pi$, $\eta_{0}$, $\eta_{1}$ (for instance, using the sample splitting techniques described in Section 4.2.3 below), which shall be denoted by $\hat{\pi}$, $\hat{\eta}_{0}$ and $\hat{\eta}_{1}$, respectively. Then, one may calculate pseudo-observations as
@@
for $i\in\{1,\ldots,n\}$. With these pseudo-observations in place, we may apply the methods described in the previous subsections to the data $(X_{1},\tilde{Y}_{1}),\ldots,(X_{n},\tilde{Y}_{n})$. While this approach allows for a high degree of flexibility and to leverage the predictive capabilities of modern machine learning tools, finite-sample guarantees are difficult to achieve. Section 4.2 below investigates the performance empirically in a simulation study, where the data-generating distributions are chosen to closely match those observed for real-world clinical data using software by sun2024benchtm.
@@ 4 Simulation study
The work by wan2024confidence and mueller2024isotonic contains proofs of the Type I error guarantee under suitable assumptions for the main methods presented in Section 3. The aim of this simulation study is to identify how violations of these assumptions affect the ability of the methods to correctly identify subgroups while controlling Type I error rates. More specifically, we are interested in clinical trial settings like those presented in Section 2. While in such settings one may have some domain knowledge of the data-generating mechanism, one will sometimes not be able to say with certainty that the assumptions of either presented method are fully satisfied. Hence, before we can use our methods to identify subgroups of high risk of adverse events (AEs) (Section 2.1) or high treatment effect with our study data (Section 2.2), we need to understand the reliability of these results when the true mechanisms do not align with the assumptions needed to guarantee Type I error rate control. Moreover, we also seek to quantify the effectiveness of the discussed methods at selecting subgroups of interest.
@@
We now first present the simulation study for the application of identifying subgroups at high risk of AEs, before continuing to the problem of treatment effect heterogeneity (TEH). For both studies, the presentation of setup and results will follow the ADEMP (Aims, Data, Estimand, Methods and Population) framework (morris2019using).
@@ 4.1.1 Data generation
Firstly, we sample covariate values that mimic the covariate distribution of the original case study in Section 2.1 by using the package syn-function from the R-package synthpop (nowok2016synthpop) with default parameters. Secondly, at each sampled covariate point $X_{i}$, we independently sample the response $Y_{i}$ from a Bernoulli distribution with success probability $\eta(X_{i})$ for the chosen ground truth $\eta$.
@@ 4.1.3 Methods
For the GLM-based method in Section 3.1, one needs to specify a GLM (i.e. which covariates to include, whether to include interaction effects, a link function, etc.In each application of the GLM-based method, the critical values $\hat{c}_{1}$ and $\hat{c}_{2}$ appearing in Section 3.1 have to be found via simulation, whose number of replications we also need to specify. Throughout the simulation study at hand, we used $1000$ replications per application of the method. We refer the reader to wan2024confidence and references therein for details on how $\hat{c}_{1}$ and $\hat{c}_{2}$ are found.). Throughout, we use all covariates as predictors and do not include any interaction effects. Moreover, noting that we are in a classification setting as our responses are binary, we use a logit link function.
@@ Results
In Section 3.2.3, we present the antichaining approach. Figure 6 illustrates that this approach can protect against the large Type I error rates incurred due to directional misspecification of the binary variable, without significant deviation from the power achieved by an application of ISS with correct directional specification. Indeed, it can even lead to an increase in power, as the problem of scaling the binary variable appropriately disappears. As such, the antichaining method is a useful addition to the ISS method as presented by mueller2024isotonic and extends its applicability even further.
@@ 4.2.1 Data generation
We use the benchtm package in R (sun2024benchtm) to generate simulated datasets. This package allows for the creation of data with varying degrees of treatment effect heterogeneity, designed to mimic real clinical trial data. The distributions of the simulated variables are based on those from a Phase III study of a compound targeting an inflammatory disease (see sun2024benchtm for more details).
@@ 4.2.3 Methods
We now give details on the high-level description of the two-step learner method described in Section 3.3. Here, we let $X_{i}$ denote the $i$th covariate vector, where we only include relevant variables, i.e. all those that are either prognostic or predictive.
@@ 5.2 Revisiting case study 2 (efficacy): Subgroup selection for desired treatment response
Subgroups were defined using two continuous variables that had been identified as promising effect modifiers in previous studies (Sechidis2021; bornkamp2023predicting). As noted earlier, the goal of these case studies is to demonstrate the utility of the considered methods, rather than to make claims about treatment efficacy. Importantly, because our analysis is post-hoc and the variables included were identified as promising effect modifiers from previous works that used the same data, it is not reasonable to expect any formal control of the Type I error rate. We refrain from using the original names of these variables and will refer to them as $X^{(1)}$ and $X^{(2)}$, scaling their values to lie within the interval $[0,1]$. Furthermore, we know the monotonicity of the effects; the larger the $X^{(1)}$ the higher the effect, while the opposite holds for $X^{(2)}$, the lower the higher the effect.
The estimand of interest in our example is the mean difference, which can be expressed as: $E(Y\mid T=1,X^{(1)},X^{(2)})-E(Y\mid T=0,X^{(1)},X^{(2)})\geq\tau,$ with threshold $\tau=10$ that is selected for illustrative purposes to represent a high treatment effect on the mean difference scale. This implies that we want to identify the subgroup of patients where the expected response on the patients that take the experimental drug is at least 10 units higher than those receiving placebo. To ensure statistical rigor, we controlled the Type I error rate at a significance level of $\alpha=0.10,$ a level widely used for subgroup analysis (dane2019subgroup).
@@ 6.2 Discussion on practical considerations
Our work aligns with ongoing regulatory discussions. For instance, the recent FDA draft guidance (FDA2025) on the use of artificial intelligence in regulatory decision making explicitly highlights subgroup identification in safety monitoring as a key motivating use case. By offering methods that combine flexibility with strong statistical guarantees, our framework addresses a critical regulatory need: enabling data-driven discovery while ensuring reliability and reproducibility. In safety focused applications, these guarantees help regulators build confidence that identified subgroups genuinely meet predefined safety thresholds, thereby reducing unnecessary monitoring without compromising patient protection. In efficacy focused applications, such guarantees can support claims of differential benefit across populations, an increasingly important consideration in labeling decisions and precision medicine initiatives.
@@
Finally, for controlled subgroup selection methods to have meaningful impact in drug development, they must be embedded within structured, pre-planned workflows. The WATCH framework (Workflow to Assess Treatment Effect Heterogeneity) exemplifies such an approach, providing a systematic process that begins with prespecified analysis planning (sechidis2025watchworkflowassesstreatment). Once initial data analysis and preprocessing are complete, the methods presented in this paper can be applied to perform controlled subgroup selection in a principled and reproducible manner.
@@ Acknowledgments
The authors would like to thank Alain Marti for his feedback on the manuscript.

### v2

@@ 3.3 Subgroup selection for heterogeneous treatment effects
Now, if in addition to the observed response $Y_{i}(T_{i})$, we would observe the counterfactual response $Y_{i}(1-T_{i})$, we could take $Y_{i}(1)-Y_{i}(0)$ as the input for the response in the methods described in the previous subsections, as the covariate-conditional expectation of that difference is exactly the CATE, so that the output of the above methods would give an estimate of $S_{\tau}$. Unfortunately, we do not observe the counterfactual response, and this is where the meta-learning framework may be useful.
A range of meta-learning techniques have been discussed in the literature (see for instance the review by Lipkovich et al. (2024)) and we will be focusing on using the so-called doubly robust (DR) learner for generating pseudo-observations mimicking the ITEs (Kennedy, 2023). The idea here is to use flexible machine learning methods in order to estimate the propensity $\pi:x\mapsto\mathbb{P}(T_{i}=1\mid X_{i}=x)$ and the treatment-conditional regression functions $\eta_{t}:x\mapsto\mathbb{E}\bigl\{Y_{i}(T_{i})\mid T_{i}=t,X_{i}=x\bigr\}$ for $t\in\{0,1\}$. Writing $\hat{\pi}$ and $\hat{\eta}_{t}$ for estimators of these quantities, we then generate pseudo-observations as
@@
for $i\in\{1,\ldots,n\}$ using sample splitting (as will be described in more detail in Section 4.2.3). The construction of the pseudo-observations $\tilde{Y}_{i}$, $i\in\{1,\ldots,n\}$, as proxies for the individual treatment effects $\mathrm{ITE}_{i}$, $i\in\{1,\ldots,n\}$, in this way is rooted in the theory of efficient influence functions for estimating the CATE when neither the outcome functions $\eta_{t}$, $t\in\{0,1\}$, nor the propensity $\pi$ are known (Hines et al., 2022). Indeed, even though we are interested in the analysis of randomised control trials, where the propensity is known, we nonetheless focus on the DR learner due to its robustness against misspecification of the outcome model as observed by Sechidis et al. (2025b).
With these pseudo-observations in place, we may apply the methods described in the previous subsections to the data $(X_{1},\tilde{Y}_{1}),\ldots,(X_{n},\tilde{Y}_{n})$. While this approach allows for a high degree of flexibility and to leverage the predictive capabilities of modern machine learning tools, finite-sample guarantees are difficult to achieve. Section 4.2 below investigates the performance empirically in a simulation study, where the data-generating distributions are chosen to closely match those observed for real-world clinical data using software by Sun et al. (2024).
@@ 4 Simulation study
The work by Wan et al. (2024) and Müller et al. (2025) contains proofs of the Type I error guarantee under suitable assumptions for the main methods presented in Section 3. The aim of this simulation study is to identify how violations of these assumptions affect the ability of the methods to correctly identify subgroups while controlling Type I error rates. More specifically, we are interested in clinical trial settings like those presented in Section 2. While in such settings one may have some domain knowledge of the data-generating mechanism, one will sometimes not be able to say with certainty that the assumptions of either presented method are fully satisfied. Hence, before we can use our methods to identify subgroups of high risk of adverse events (AEs) (Section 2.1) or high treatment effect with our study data (Section 2.2), we need to understand the reliability of these results when the true mechanisms do not align with the assumptions needed to guarantee Type I error rate control. Moreover, we also seek to quantify the effectiveness of the discussed methods at selecting subgroups of interest.
@@
We now first present the simulation study for the application of identifying subgroups at high risk of AEs, before continuing to the problem of treatment effect heterogeneity (TEH). For both studies, the presentation of setup and results will follow the ADEMP (Aims, Data, Estimand, Methods and Population) framework (Morris et al., 2019).
@@ 4.1.1 Data generation
Firstly, we sample covariate values that mimic the covariate distribution of the original case study in Section 2.1 by using the package syn-function from the R-package synthpop (Nowok et al., 2016) with default parameters. Secondly, at each sampled covariate point $X_{i}$, we independently sample the response $Y_{i}$ from a Bernoulli distribution with success probability $\eta(X_{i})$ for the chosen ground truth $\eta$.
@@ 4.1.3 Methods
For the GLM-based method in Section 3.1, one needs to specify a GLM (i.e. which covariates to include, whether to include interaction effects, a link function, etc.In each application of the GLM-based method, the critical values $\hat{c}_{1}$ and $\hat{c}_{2}$ appearing in Section 3.1 have to be found via simulation, whose number of replications we also need to specify. Throughout the simulation study at hand, we used $1000$ replications per application of the method. We refer the reader to Wan et al. (2024) and references therein for details on how $\hat{c}_{1}$ and $\hat{c}_{2}$ are found.). Throughout, we use all covariates as predictors and do not include any interaction effects. Moreover, noting that we are in a classification setting as our responses are binary, we use a logit link function.
@@ Results
In Section 3.2.3, we present the antichaining approach. Figure 6 illustrates that this approach can protect against the large Type I error rates incurred due to directional misspecification of the binary variable, without significant deviation from the power achieved by an application of ISS with correct directional specification. Indeed, it can even lead to an increase in power, as the problem of scaling the binary variable appropriately disappears. As such, the antichaining method is a useful addition to the ISS method as presented by Müller et al. (2025) and extends its applicability even further.
@@ 4.2.1 Data generation
We use the benchtm package in R (Sun et al., 2024) to generate simulated datasets. This package allows for the creation of data with varying degrees of treatment effect heterogeneity, designed to mimic real clinical trial data. The distributions of the simulated variables are based on those from a Phase III study of a compound targeting an inflammatory disease (see Sun et al. (2024) for more details).
@@ 4.2.3 Methods
We now give details on the high-level description of the doubly robust learner method described in Section 3.3. Here, we let $X_{i}$ denote the $i$th covariate vector, where we only include relevant variables, i.e. all those that are either prognostic or predictive.
@@ 5.2 Revisiting case study 2 (efficacy): Subgroup selection for desired treatment response
Subgroups were defined using two continuous variables that had been identified as promising effect modifiers in previous studies (Sechidis et al., 2021; Bornkamp et al., 2024). As noted earlier, the goal of these case studies is to demonstrate the utility of the considered methods, rather than to make claims about treatment efficacy. Importantly, because our analysis is post-hoc and the variables included were identified as promising effect modifiers from previous works that used the same data, it is not reasonable to expect any formal control of the Type I error rate. We refrain from using the original names of these variables and will refer to them as $X^{(1)}$ and $X^{(2)}$, scaling their values to lie within the interval $[0,1]$. Furthermore, we know the monotonicity of the effects; the larger the $X^{(1)}$ the higher the effect, while the opposite holds for $X^{(2)}$, the lower the higher the effect.
The estimand of interest in our example is the mean difference, which can be expressed as: $E(Y\mid T=1,X^{(1)},X^{(2)})-E(Y\mid T=0,X^{(1)},X^{(2)})\geq\tau,$ with threshold $\tau=10$ that is selected for illustrative purposes to represent a high treatment effect on the mean difference scale. This implies that we want to identify the subgroup of patients where the expected response on the patients that take the experimental drug is at least 10 units higher than those receiving placebo. To ensure statistical rigor, we controlled the Type I error rate at a significance level of $\alpha=0.10,$ a level widely used for subgroup analysis (Dane et al., 2019).
@@ 6.2 Discussion on practical considerations
Our work aligns with ongoing regulatory discussions. For instance, the recent FDA draft guidance (U.S. Food and Drug Administration, 2025) on the use of artificial intelligence in regulatory decision making explicitly highlights subgroup identification in safety monitoring as a key motivating use case. By offering methods that combine flexibility with strong statistical guarantees, our framework addresses a critical regulatory need: enabling data-driven discovery while ensuring reliability and reproducibility. In safety focused applications, these guarantees help regulators build confidence that identified subgroups genuinely meet predefined safety thresholds, thereby reducing unnecessary monitoring without compromising patient protection. In efficacy focused applications, such guarantees can support claims of differential benefit across populations, an increasingly important consideration in labeling decisions and precision medicine initiatives.
@@
Finally, for controlled subgroup selection methods to have meaningful impact in drug development, they must be embedded within structured, pre-planned workflows. The WATCH framework (Workflow to Assess Treatment Effect Heterogeneity) exemplifies such an approach, providing a systematic process that begins with prespecified analysis planning (Sechidis et al., 2025a). Once initial data analysis and preprocessing are complete, the methods presented in this paper can be applied to perform controlled subgroup selection in a principled and reproducible manner.
@@ Acknowledgments
The authors would like to thank Alain Marti, Jelle J. Goeman and Rajen D. Shah for their feedback on the manuscript.